# DiG-Net: Enhancing Quality of Life through Hyper-Range Dynamic Gesture Recognition in Assistive Robotics

### v1

@@ 3.4 DiG-Net analysis
Figure 7: Gesture recognition success rate of the DiG-Net model with regard to the number of labeled training images.
The above results were acquired with a video window length of up to $n=84$ frames. Hence, we now evaluate the performance of the DiG-Net with respect to the window length. The DiG-Net model was trained multiple times on video sequences $V_{t}$ of varying lengths $n$, and each trained model’s performance was evaluated on test sequences of similar lengths. Figure 8 shows the recognition success rate of the DiG-Net with regard to the number of frames $n$ in a video sample. First, with only one image of the gesture (i.e., $n=1$), a poor recognition success rate of $72.5\%$ is achieved, due to the lack of dynamic information. As more frames are added to $V_{t}$, more dynamic information is included and the recognition improves. A longer video encapsulates more dynamic information on the exhibited gesture, yielding better recognition accuracy. However, increasing $n$ may affect the time sensitivity and frequency of the recognition in real time. Nevertheless, increasing the sequence length beyond $n=84$ frames offers diminishing returns in terms of accuracy.
Figure 8: Gesture recognition success rate of the DiG-Net concerning the number of frames $n$ in a video.
Table 4: Ablation study results for DiG-Net components
Model Variant Success rate (%) DWA GSS $F_{1}$ score mAP (%)
DiG-Net w/o DADA module 88.9 0.84 0.86 0.87 85.2
DiG-Net w/o STG module 89.7 0.85 0.88 0.88 86.4
DiG-Net w/o Graph Transformer 87.5 0.83 0.85 0.86 84.0
DiG-Net w/o RSTDAL 90.1 0.86 0.89 0.89 87.1
DiG-Net with short sequence 91.2 0.87 0.90 0.91 89.3
Full DiG-Net model 97.3 0.92 0.96 0.93 94.9
@@
Figure 9: Success rate of the DiG-Net model after fine-tuning to recognize five additional gestures, plotted against the number of video clips used for fine-tuning (each clip spans 84 frames).
@@
3.5 Sequence Analysis and Varying Conditions
@@
Table 5: Results of Sequence Analysis and Varying Conditions
Metric Success rate (%)
Sequence accuracy 94.4%
Controlled lighting (Indoor) 96.8%
Bright sunlight (Outdoor) 93.1%
Overcast (Outdoor) 91.5%
3.6 Robustness and Inference Efficiency
@@
Table 6: Recognition Accuracy under Environmental Noise Conditions
Condition Success rate (%)
Mild clutter/interference 95.8%
Moderate clutter/interference 93.5%
Severe clutter/interference 90.1%
@@
Table 7: Inference Speed under Real-Time Simulation
Sequence Length Inference Speed (FPS)
@@
Table 8: Recognition Accuracy under Optical Degradation
Degradation Level Success rate (%)
Mild blur/fog 94.7%
Moderate blur/fog 91.2%
Severe blur/fog 88.3%
4 Conclusions
@@
The model significantly enhances the quality of life and independence for individuals with mobility impairments by enabling intuitive, long-range interaction with robotic assistants across various environments, including home healthcare, industrial safety, and emergency response. Future work will aim at expanding the assistive gesture vocabulary, integrating non-intentional body language cues, improving real-time responsiveness, and enhancing robustness to highly dynamic and crowded environments, thus further broadening the impact and usability of assistive robotic technologies.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 3.4. DiG-Net analysis
Figure 7. Gesture recognition success rate of the DiG-Net model with regard to the number of labeled training images.
A line plot showing the recognition success rate of the DiG-Net model versus the number of labeled training images. The curve rises sharply as more data are added and gradually plateaus, indicating that model accuracy saturates beyond roughly 25,000 labeled samples. Axes are labeled with the number of training images on the x-axis and success rate on the y-axis.
The above results were acquired with a video window length of up to $n=84$ frames. Hence, we now evaluate the performance of the DiG-Net with respect to the window length. The DiG-Net model was trained multiple times on video sequences $V_{t}$ of varying lengths $n$, and each trained model’s performance was evaluated on test sequences of similar lengths. Figure 8 shows the recognition success rate of the DiG-Net with regard to the number of frames $n$ in a video sample. First, with only one image of the gesture (i.e., $n=1$), a poor recognition success rate of $72.5\%$ is achieved, due to the lack of dynamic information. As more frames are added to $V_{t}$, more dynamic information is included, and the recognition improves. A longer video encapsulates more dynamic information on the exhibited gesture, yielding better recognition accuracy. However, increasing $n$ may affect the time sensitivity and frequency of the recognition in real time. Nevertheless, increasing the sequence length beyond $n=84$ frames offers diminishing returns in terms of accuracy.
Figure 8. Gesture recognition success rate of the DiG-Net concerning the number of frames $n$ in a video.
A line chart showing the gesture recognition success rate of the DiG-Net model as a function of the number of frames per video sequence. The x-axis represents the number of frames (\$n\$), and the y-axis shows recognition success rate. The curve rises steadily as more frames are used, indicating improved temporal modeling, and plateaus once the sequence length exceeds around 40 frames.
@@
Table 4. Ablation study results for components of the DiG-Net model. Each variant excludes one key module to evaluate its contribution to overall performance.
[…]
Figure 9. Success rate of the DiG-Net model after fine-tuning to recognize five additional gestures, plotted against the number of video clips used for fine-tuning (each clip spans 84 frames).
@@
3.5. Human–Robot Evaluation and Comparative Insights
To better contextualize the technical results of DiG-Net, we compared its quantitative performance with a user-based evaluation examining how humans perceive gestures at similar long-range conditions. This comparative perspective bridges model evaluation with human interpretability, an aspect central to HRI. The results of the user study (Section 2.3) are summarized in Table 5. Ten participants (3 female, 7 male; age range 26–74; all with normal or corrected vision and no prior robotics experience) took part in the study. Participants rated their confidence on a 7-point Likert scale after each trial (1 = very uncertain, 7 = very confident), reflecting how sure they were about correctly identifying the gesture. Recognition accuracy and confidence decreased with distance, particularly for static gestures, whereas dynamic gestures remained comparatively easier to identify. Participants reported that temporal motion provided essential cues for understanding intent when spatial detail was limited. These findings mirror DiG-Net’s design philosophy, which explicitly leverages motion-aware and depth-compensated features to address similar perceptual degradations.
Table 5. Summary of human recognition performance across gesture types and distances.
Condition Accuracy (%) $\uparrow$ Reaction Time (s) $\downarrow$ Confidence (1–7) $\uparrow$
Static (Short Range) 96.0 1.2 6.5
Static (Mid Range) 82.0 1.9 5.3
Static (Long Range) 68.0 2.5 4.7
Dynamic (Short Range) 98.0 1.1 6.7
Dynamic (Mid Range) 91.0 1.6 6.2
Dynamic (Long Range) 84.0 1.8 5.8
Values in Table 5 represent mean performance across all participants for each condition. While the study was exploratory and limited in scale, the reported values reflect consistent trends across participants. When comparing these behavioral findings to the model’s quantitative results, we observe that DiG-Net consistently outperforms human recognition at extended ranges, especially for static gestures where visual detail diminishes. For example, while participants achieved an average accuracy of 84% for dynamic gestures at a long distance, DiG-Net maintained 94.9% under the same conditions. This demonstrates that the model not only mirrors human motion sensitivity but also compensates for perceptual limitations that emerge with increasing range. Despite the modest sample size, the observed trends align with prior HRI perception studies and offer valuable qualitative insight into the interpretability of long-range gestures.
3.6. Sequence Analysis and Varying Conditions
@@
Table 6. Results of Sequence Analysis and Varying Conditions
Metric Success rate (%) $\uparrow$
Sequence accuracy 94.4
Controlled lighting (indoor) 96.8
Bright sunlight (outdoor) 93.1
Overcast (outdoor) 91.5
3.7. Robustness and Inference Efficiency
@@
Table 7. Recognition Accuracy under Environmental Noise Conditions
Condition Success rate (%) $\uparrow$
Mild clutter/interference 95.8
Moderate clutter/interference 93.5
Severe clutter/interference 90.1
@@
Table 8. Inference Speed under Real-Time Simulation
Sequence length (frames) Inference speed (FPS) $\downarrow$
@@
Deployment Considerations: To evaluate DiG-Net’s suitability for practical assistive robotics deployment, we conducted real-time experiments on an NVIDIA Jetson Orin Nano embedded platform. The model processed gesture sequences of 8–32 frames at 15–25 frames per second, consistent with the simulated inference rates reported in Table 8. These results were achieved using the full-precision (FP32) model without any compression, quantization, pruning, or TensorRT acceleration, operating entirely within the platform’s 8 GB memory limit. This confirms that DiG-Net supports real-time operation on resource-constrained embedded hardware commonly used in mobile robotics applications.
Table 9. Recognition Accuracy under Optical Degradation
Degradation level Success rate (%) $\uparrow$
Mild blur/fog 94.7
Moderate blur/fog 91.2
Severe blur/fog 88.3
4. Conclusions
@@
The model significantly enhances the quality of life and independence for individuals with mobility impairments by enabling intuitive, long-range interaction with robotic assistants across various environments, including home healthcare, industrial safety, and emergency response. Future work will focus on expanding the assistive gesture vocabulary, integrating non-intentional body language cues, improving real-time responsiveness, and enhancing robustness in dynamic and crowded environments. We acknowledge that the user study sample size (N=10) is limited. Future work should include larger and more demographically diverse participant pools to strengthen the generalizability of the HRI findings. These directions aim to bridge perceptual modeling with human-centered interaction design, contributing to the development of more transparent, adaptive, and socially aware assistive robotic systems.
In addition to the above contributions, we recognize several limitations that outline directions for future work. Our gesture dataset was collected from participants recruited in a single geographic region, and ethnicity and hand-size measurements were not explicitly balanced or recorded. In the accompanying user study, participants spanned a wider age range (26–74 years). At hyper-range distances, fine-grained hand morphology is strongly attenuated by resolution and imaging effects; accordingly, DiG-Net is designed to rely primarily on coarse spatiotemporal motion cues and full-body context rather than detailed hand-shape cues. Nevertheless, future work will aim to expand the dataset to include greater ethnic diversity to ensure broad generalization.
Ethical Considerations
All user study procedures followed ACM’s ethical guidelines for research involving human participants. All participants provided informed consent before participation, and all data were anonymized before analysis. The study did not involve any automated decision-making affecting individuals, and no personally identifiable information was collected or stored.
Conflict of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this paper.
@@
Acknowledgments
The authors thank the participants and colleagues who assisted in the user study and provided valuable feedback.